OpenStudy (anonymous):

Can someone please me to solve this: 3√-64

OpenStudy (anonymous):

is the sqrt or cubed root?

OpenStudy (anonymous):

square root

OpenStudy (anonymous):

\[3\sqrt{-64} = 24i\]

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

if it's cubed , then[ (-2^3)^2 ](1/3)=4

OpenStudy (anonymous):

using radicals

OpenStudy (anonymous):

if it sqrt , 3*sqrt(64) i = 24 i

OpenStudy (anonymous):

\[3\sqrt{-64} = 3\sqrt{-1}\sqrt{64} = 3*i*8 = 24i\]

OpenStudy (anonymous):

what i stands for?

OpenStudy (anonymous):

I assume you're confused about the question, if you don't know what i is then you probably mean: \[\sqrt[3]{-64}\] In which case, the answer is -4 (-4)^3 = -64

OpenStudy (anonymous):

i am going to make a bet that it says \[\sqrt[3]{-64}\] so it is \[\sqrt[3]{-64}=-4\] because \[(-4)^3=-4\times -4\times -4=-64\]

OpenStudy (anonymous):

@alfers read "cubed root of -64" it means the number whose cube is -64

OpenStudy (anonymous):

the little 3 is called the "index" and if you don't see one that means square root but for example \[\sqrt[4]{16}=2\] because \[2^4=16\]

OpenStudy (anonymous):

oh okay i get it. thank u so much how about this one? 5√-32 ??

OpenStudy (anonymous):

think of a number that when you raise it to the fifth power you get -32

OpenStudy (anonymous):

i give you a hint, it can't be very big because it is going to be -32 when you raise it to the fifth power also it must be negative

OpenStudy (anonymous):

now i get it thank u so much!